A pharmacist mixed some 10%-saline solution with some 15%-saline solution to obtain 100 ml of a 12%-saline solution. How much of the 10%-saline solution did the pharmacist use in the mixture?
Answer by rwm(914)

.1*x+.15*y=.12*100
x+y=100
x=60 y=40
check
.1*60+.15*40=.12*100